# Prospect of Precision Cosmology and Testing General Relativity using Binary Black Holes- Galaxies Cross-correlation

### v1

@@ 1 Introduction
The General Theory of Relativity (GR) has been a foundational element in our understanding of gravity for over a century, offering an elegant description of how mass interacts with spacetime curvature. While GR has proven remarkably successful, it encounters challenges under extreme conditions such as those near black holes and during the early universe (Thorne, 1995; Sathyaprakash & Schutz, 2009). Recently, a significant breakthrough occurred with the direct detection of gravitational waves (GW) by the LIGO-Virgo collaboration (Abbott et al., 2016b, c; Scientific et al., 2017; Abbott et al., 2017a; Goldstein et al., 2017). This milestone has opened new pathways for testing GR, especially within the context of compact binary systems in relativistic regimes. These systems, including Binary Neutron Stars (BNS), Neutron Star-Black Hole pairs (NSBH), and Binary Black Holes (BBH), offer unique opportunities to study gravity’s fundamental nature across various mass scales and cosmic distances. Ground-based detectors like LIGO (Aasi et al., 2015), Virgo (Acernese et al., 2014), and KAGRA (Akutsu et al., 2021), operating in the hertz to kilohertz frequency range, have been crucial in this endeavor. The forthcoming addition of space-based detectors such as LISA (Amaro-Seoane et al., 2017), which will function in the millihertz frequency range, alongside advanced ground-based observatories like LIGO-Aundha (LIGO India)(Saleem et al., 2021) [^1] https://dcc.ligo.org/LIGO-M1100296/public, Cosmic Explorer (CE) (Reitze et al., 2019), and the Einstein Telescope (ET) (Punturo et al., 2010), promises to significantly enhance our capacity to test GR’s predictions and deepen our grasp of gravitational phenomena. By leveraging these combined observational capabilities, we aim to refine and expand our understanding of the intricate relationship between gravity and the cosmos. While various alternative models have been proposed to address specific scenarios, this study focuses on testing GR through a model-independent, data-driven approach to GW propagation. Through this rigorous analysis, we seek to explore and validate the fundamental principles of GR across a wide range of conditions and scales. GW and EM signals’ propagation speed and luminosity distance offer insight into modified gravity theories. Factors like graviton mass, frictional term(due to running effect of Planck mass) and anisotropic source term contribute to deviations (Deffayet & Menou, 2007; Saltas et al., 2014; Nishizawa, 2018a; Belgacem et al., 2018a, b; Lombriser & Taylor, 2016; Lombriser & Lima, 2017).
@@
Various model-dependent parameterizations of $\rm{\gamma(z)}$ have been proposed in the literature. One such parameterization is the $\rm{(\Xi_{0},n)}$ model (Belgacem et al., 2018b), which has been shown to be detectable with LVK (Leyde et al., 2022) and LISA (Baker & Harrison, 2021). This parameterization relies on two positive parameters, $\rm{\Xi_{0}}$ and $\rm{n}$, where setting $\rm{\Xi_{0}=1}$ corresponds to General Relativity (GR). Numerous modified gravity models predict the frictional term under different scenarios. A prominent class among these is Scalar-Tensor theories, which include a scalar degree of freedom crucial for the universe’s evolution. Notable scalar-tensor theories include Brans-Dicke theory (Brans & Dicke, 1961), $f(R)$ gravity (Hu & Sawicki, 2007), and covariant Galileon models (Chow & Khoury, 2009), all of which fall under Horndeski theories. Beyond Horndeski theories, Degenerate Higher Order Scalar-Tensor (DHOST) theories (Frusciante et al., 2019; Crisostomi et al., 2016; Achour et al., 2016) have been developed as the most general scalar-tensor theories propagating a single scalar degree of freedom along with the helicity-2 mode of a massless graviton.
@@
To address these challenges, it is essential to test GR in a model-independent manner and measure the integrated effect of $\rm{\gamma(z)}$ as a function of the data, introduced for siren in our previous analysis (Afroz & Mukherjee, 2023, 2024). This function $\rm{\mathcal{F}(z)}$ is designed to capture any deviation from GR and serves as a metric to assess the difference between the distances $\mathrm{D_{l}^{GW}(z)}$ and $\mathrm{D_{l}^{EM}(z)}$. In GR, $\rm{\mathcal{F}(z)=1}$ at all redshifts, and our approach makes it possible to reconstruct this in a redshift tomographic-way. In this framework, along with $\rm{\mathcal{F}(z)}$, we can also infer the Hubble constant, $\rm{H_{0}}$ making it possible for jointly explore cosmology and modified gravity theories.
In our previous work, using bright standard sirens (GW sources with EM counterpart), we demonstrated a model-independent reconstruction of the frictional term for various sources, utilizing both current and upcoming ground-based detectors (Afroz & Mukherjee, 2023, 2024). However, our previous analysis was limited to BNS and NSBH systems, as these are the most likely to have EM counterparts. Detecting EM counterparts in BNS and NSBH mergers presents significant challenges due to factors such as prompt collapse, ejecta mass, and orientation. The detectability of collimated jets depends on their opening angles, with wider angles increasing visibility. Additionally, detection is complicated by distance, instrument sensitivity, and rapid event localization, with phenomena such as "structured jets" and "off-axis viewing" further impacting observations (Rees & Mészáros, 1998; Ramirez-Ruiz et al., 2001; Kumar & Piran, 2000). For the majority of BNS and NSBH events, EM counterparts may not be detectable, especially if baryonic matter is absent in the environments of stellar-mass binary black holes. Notably, the most common class of binary mergers, BBHs, are unlikely to exhibit EM counterparts. Consequently, like other bright siren analyses, our earlier study was confined to sources from which EM counterparts are anticipated, such as BNS and NSBH mergers, provided there is dedicated EM follow-up. To address these limitations, in this study, we have adopted a method that facilitates the exploration of alternative theories of gravity using redshift unknown GW sources. This technique exploits the three-dimensional spatial clustering of GW sources with galaxy redshift surveys to determine their host redshifts through cross-correlation(Mukherjee & Wandelt, 2018; Mukherjee et al., 2020, 2021a; Bera et al., 2020; Mukherjee et al., 2021b; Oguri, 2016; Diaz & Mukherjee, 2022; Ghosh et al., 2023)
@@ 2.1 Data-Driven Reconstruction of F(z)
The comoving sound horizon $\rm{r_{s}}$ at the drag epoch can be related to the comoving sound horizon $\rm{r_{*}}$ at the redshift of recombination ($\rm{z_{*}=1090}$) by $\rm{r_{d}\approx 1.02r_{*}}$. The quantity $r_{*}$ is inferred from CMB observation, using the position of the first peak in the CMB temperature power spectrum $\rm{\theta_{*}=r_{*}/(1+z_{*})D^{EM}_{A}(z_{*})}$ (Spergel et al., 2007; Hinshaw et al., 2013; Aghanim et al., 2020). As a result, the product of all these observable quantities will lead to an identity value at all redshifts if GR and the fiducial $\Lambda$CDM model of cosmology is the correct theory. However, if there is any departure from these models, then we can measure a deviation from one with redshift. It is important to note that the value of $\rm{r_{s}}$ depends on the redshift above $\rm{z_{d}=1020}$. Though it depends on the model of cosmology, it is a redshift independent number of a constant value for the low redshifts $\rm{z<2}$. So, even if there is an inaccuracy in the inference of true $\rm{r_{s}}$, the value of $\rm{\mathcal{F}(z)}$ will vary by an overall normalization. But the reconstructed redshift evolution of $\rm{\mathcal{F}(z)}$ will remain the same. In one of the later sections, we will show results for both the cases, (i) only $\rm{\mathcal{F}(z)}$ and (ii) $\rm{\mathcal{F}(z)}$ and $\rm{H_{0}}$ together. The second case will make it possible to marginalise over any inaccurate inference of sound horizon $\rm{r_{s}}$ due to incorrect inference of the Hubble constant. In the above equation, $\rm{D_{l}^{GW}(z)}$ is measured from GW sources, $\rm{\theta_{BAO}}$ is measured from galaxy two point correlation function, and $\rm{r_{s}}$ is inferred from the CMB temperature fluctuations. The redshift of the source can be inferred using cross-correlation of galaxy surveys with dark sirens. In Subsection 4, we provide a detailed framework for inferring redshift using cross-correlation with redshift-known galaxies.
@@ 3.1 Modelling of the astrophysical population of GW source
Merger rate: In this study we use a delay time model of the binary mergers (O’shaughnessy et al., 2010; Dominik et al., 2015). In the delay time model, the merger rate is described in terms of the delay time distribution, denoted as $\rm{t_{d}}$. The delay time refers to the elapsed time between the formation of stars that will eventually become black holes and the actual merging of these black holes. It is important to note that the time delay is not uniform across all binary black holes but instead follows a specific distribution. This distribution function accounts for the variations in the delay time and is defined as follows
@@
where $T_{obs}$ indicates the total observation time, $\rm{\frac{dV_{c}}{dz}}$ corresponds to the comoving volume element, and $\rm{R(z)}$ denotes the merger rate (Karathanasis et al., 2022). We consider the delay time merger rate with a specific minimum delay time $\rm{t_{d}=500\text{Myrs}}$ and a power-law exponent of $\rm{d=1}$. To determine which events are detectable, the calculation of the matched filtering signal-to-noise ratio (SNR) plays a crucial role. The SNR serves as a measure of the strength of the GW signal relative to the background noise. Only those events with a matched filtering SNR greater than or equal to a predetermined threshold SNR ($\rm{\rho_{\rm TH}}$) can be reliably detected. (Maggiore, 2007).
@@
In this study, we employ two sets of GW detectors: one comprising LIGO (Aasi et al., 2015), Virgo (Acernese et al., 2014), and KAGRA (Akutsu et al., 2021) (LVK), and the other consisting of CE with two different configurations: a 20 km configuration (CE(20Km)+ET) and a 40 km configuration (CE(40Km)+ET), along with ET (Punturo et al., 2010). We establish different SNR thresholds for different GW observatories: $\rm{\rho_{\rm Th}=12}$ for the LVK (O5 sensitivity) system and $\rm{\rho_{\rm Th}=20}$ for both configurations of the CE+ET. We adopt a redshift bin size of $\rm{\Delta z=0.05}$, within which the total number of events is computed using Equation 9. Event generation is based on distances derived from redshifts, parameter $\rm{\Theta}$, and the mass distributions of the objects. Over a five-year operational period with a 75% duty cycle, the expected number of detectable events is projected to be approximately $\sim$13000 for LVK, $\sim$177000 for CE(40Km)+ET, and $\sim$144000 for CE(20Km)+ET, assuming a local merger rate ($\rm{R_{0}}$) of 20 $\rm{Gpc^{-3}yr^{-1}}$. It is important to note that for this analysis, we have specifically utilized the CE(20 km)+ET configuration. From now onwards, this combination is defined as CEET.
@@
In Figure 2, we present the projected total and detectable BBH events within specific redshift intervals ($\rm{\Delta z=0.05}$). To obtain these estimates, we first calculate the number of BBH mergers for each redshift bin using Equation 9. For each BBH event, we use the inverse transform method to determine the mass of the two black holes and the parameter $\rm{\Theta}$, drawing from their respective probability distributions. The masses of the black holes are generated within the range of $\rm{5M_{\odot}}$ to $\rm{50M_{\odot}}$, and $\rm{\Theta}$ is considered within the interval from 0 to 4. After incorporating the redshift information for each event, we calculate the corresponding luminosity distance and the SNR using Equation 14 for a single detector. The total SNR, denoted as $\rm{\rho_{total}}$, is calculated by combining the individual SNRs from the network of detectors using the formula $\rm{\rho_{total}=\sqrt{\sum_{i}\rho_{i}^{2}}}$. This analysis considers the detection capabilities of both the CEET and the LVK network over a specified observation period of 5 years with a 75% duty cycle. The curve depicted in the graph illustrates the distribution of these events, providing insights into the temporal occurrence of these mergers through cosmic history and their detection likelihood with current and future GW observatories.
@@ 4 Cross-Correlation Technique for Redshift Estimation of Dark Sirens
where $\rm{\delta^{r}_{\text{GW}}(k,z)}$ is the original density field of GW sources without sky localization impact, and $\rm{k_{\text{eff}}(z)}$ denotes a characteristic wavenumber varying with redshift $\rm{z}$. $\rm{k_{\text{eff}}(z)}$ is the comoving scale defined as
@@
where $\rm{D_{c}(z)}$ is the comoving distance at redshift $\rm{z}$. The exponential term $\rm{e^{-\frac{k^{2}}{k_{\text{eff}}^{2}(z)}}}$ quantifies the smoothing effect caused by the sky localization error, with a rapid decrease beyond $\rm{k>k_{\text{eff}}}$. Ongoing GW detectors like LVK are helping us characterize detected GW sources. In the future, upcoming ground-based detectors such as CE, ET, and LIGO-India, along with space-based detectors like LISA, will enhance our ability to characterize these sources. These advancements will improve our measurements of the redshifts of GW sources, thereby providing better insights into the Universe’s large scale structure and evolution.
In the standard model of cosmology, galaxies trace the underlying distribution of dark matter through the galaxy bias parameter, $\rm{b_{g}(k,z)}$. Consequently, GW sources are also expected to trace the dark matter distribution, but with their own distinct bias parameter, $\rm{b_{GW}(k,z)}$, which differs from the galaxy bias parameter. The spatial clustering of galaxies and GW sources can also be described in Fourier space using the three-dimensional auto-power spectrum and cross-power spectrum at different redshifts $\rm{z}$. The matter distribution exhibits a clustering property that can be statistically described by the correlation function $\rm{\xi(r)}$, which is the fourier transform of the matter power spectrum $\rm{P_{m}(k)=\langle\delta_{m}(k)\delta_{m}(k)\rangle}$. Under the standard cosmological model, both GW events and the distribution of galaxies are biased tracers of the underlying matter distribution. Therefore, the cross-correlation of the power spectrum between GW sources and galaxy surveys can be used to infer the host redshift shell of the GW sources. For a detailed mathematical formulation, refer to Subsection 4.1.

### v2

@@ 1 Introduction
The General Theory of Relativity (GR) has been a foundational element in our understanding of gravity for over a century, offering an elegant description of how mass interacts with spacetime curvature. While GR has proven remarkably successful, it encounters challenges under extreme conditions such as those near black holes and during the early universe (Thorne, 1995; Sathyaprakash & Schutz, 2009). Recently, a significant breakthrough occurred with the direct detection of gravitational waves (GW) by the LIGO-Virgo collaboration (Abbott et al., 2016b, c; Scientific et al., 2017; Abbott et al., 2017a; Goldstein et al., 2017). This milestone has opened new pathways for testing GR, especially within the context of compact binary systems in relativistic regimes. These systems, including Binary Neutron Stars (BNS), Neutron Star-Black Hole pairs (NSBH), and Binary Black Holes (BBH), offer unique opportunities to study gravity’s fundamental nature across various mass scales and cosmic distances. Ground-based detectors like LIGO (Aasi et al., 2015), Virgo (Acernese et al., 2014), and KAGRA (Akutsu et al., 2021), operating in the hertz to kilohertz frequency range, have been crucial in this endeavor. The forthcoming addition of space-based detectors such as LISA (Amaro-Seoane et al., 2017), which will function in the millihertz frequency range, alongside advanced ground-based observatories like LIGO-Aundha (LIGO India)(Saleem et al., 2021) [^1] https://dcc.ligo.org/LIGO-M1100296/public, Cosmic Explorer (CE) (Reitze et al., 2019), and the Einstein Telescope (ET) (Punturo et al., 2010), promises to significantly enhance our capacity to test GR’s predictions and deepen our grasp of gravitational phenomena. By leveraging these combined observational capabilities, we aim to refine and expand our understanding of the intricate relationship between gravity and the cosmos. While various alternative models have been proposed to address specific scenarios, this study focuses on testing GR through a model-independent, data-driven approach to GW propagation. Through this rigorous analysis, we seek to explore and validate the fundamental principles of GR across a wide range of conditions and scales. The propagation speed of GW and EM signals, along with their corresponding luminosity distances, can provide valuable insights into modified gravity theories. Deviations from general relativity can arise due to factors such as the mass of the graviton, a frictional term ($\rm{\gamma(z)}$) associated with the running of the Planck mass, and an anisotropic source term (Deffayet & Menou, 2007; Saltas et al., 2014; Nishizawa, 2018a; Belgacem et al., 2018a, b; Lombriser & Taylor, 2016; Lombriser & Lima, 2017).
@@
Various model-dependent parameterizations of $\rm{\gamma(z)}$ have been proposed in the literature. One such parameterization is the $\rm{(\Xi_{0},n)}$ model (Belgacem et al., 2018b), which has been shown to be inferable with LVK (Mukherjee et al., 2021b; Leyde et al., 2022; Finke et al., 2021; Chen et al., 2024) and LISA (Baker & Harrison, 2021). This parameterization relies on two positive parameters, $\rm{\Xi_{0}}$ and $\rm{n}$, where setting $\rm{\Xi_{0}=1}$ corresponds to General Relativity (GR). Numerous modified gravity models predict the frictional term under different scenarios. A prominent class among these is Scalar-Tensor theories, which include a scalar degree of freedom crucial for the universe’s evolution. Notable scalar-tensor theories include Brans-Dicke theory (Brans & Dicke, 1961), $f(R)$ gravity (Hu & Sawicki, 2007), and covariant Galileon models (Chow & Khoury, 2009), all of which fall under Horndeski theories. Beyond Horndeski theories, Degenerate Higher Order Scalar-Tensor (DHOST) theories (Frusciante et al., 2019; Crisostomi et al., 2016; Achour et al., 2016) have been developed as the most general scalar-tensor theories propagating a single scalar degree of freedom along with the helicity-2 mode of a massless graviton.
@@
To address these challenges, it is essential to test GR in a model-independent manner and measure the integrated effect of $\rm{\gamma(z)}$ as a function of the data, introduced for bright siren in our previous analysis (Afroz & Mukherjee, 2023, 2024). We achieve this by using the function $\rm{\mathcal{F}(z)}$, a model-independent parameterization of $\rm{\gamma(z)}$ (more details in Section 2). This function $\rm{\mathcal{F}(z)}$ is designed to capture any deviation from GR and serves as a metric to assess the difference between the distances $\mathrm{D_{l}^{GW}(z)}$ and $\mathrm{D_{l}^{EM}(z)}$. In GR, $\rm{\mathcal{F}(z)=1}$ at all redshifts, and our approach makes it possible to reconstruct this in a redshift tomographic-way. In this framework, along with $\rm{\mathcal{F}(z)}$, we can also infer the Hubble constant, $\rm{H_{0}}$ making it possible for jointly explore cosmology and modified gravity theories.
In our previous work, using bright standard sirens (GW sources with EM counterparts), we demonstrated a model-independent reconstruction of the frictional term for various sources, utilizing both current and upcoming ground-based detectors (Afroz & Mukherjee, 2023, 2024). However, our previous analysis was limited to BNS and NSBH systems, as these are the most likely to have EM counterparts. Detecting EM counterparts in BNS and NSBH mergers presents significant challenges due to factors such as prompt collapse, ejecta mass, and orientation. While the detectability of collimated jets, such as those associated with GRBs, depends on their opening angles, wider angles increase visibility. It is important to note that kilonovae, which are also produced in these mergers, are isotropic. Thus, inclination isn’t directly responsible for the lack of redshift data from kilonovae. However, other factors like distance, instrument sensitivity, and rapid event localization still complicate the detection of EM counterparts, with phenomena such as "structured jets" and "off-axis viewing" further impacting observations (Rees & Mészáros, 1998; Ramirez-Ruiz et al., 2001; Kumar & Piran, 2000). For the majority of BNS and NSBH events, EM counterparts may not be detectable, especially if baryonic matter is absent in the environments of stellar-mass binary black holes. Notably, the most common class of binary mergers, BBHs, are unlikely to exhibit EM counterparts. Consequently, like other bright siren analyses, our earlier study was confined to sources from which EM counterparts are anticipated, such as BNS and NSBH mergers, provided there is dedicated EM follow-up. To address these limitations, in this study, we have adopted a method that facilitates the exploration of alternative theories of gravity using redshift unknown GW sources. This technique exploits the three-dimensional spatial clustering of GW sources with galaxy redshift surveys to determine their host redshifts through cross-correlation(Mukherjee & Wandelt, 2018; Mukherjee et al., 2020, 2021a; Bera et al., 2020; Mukherjee et al., 2021b; Oguri, 2016; Diaz & Mukherjee, 2022; Ghosh et al., 2023).
@@ 2.1 Data-Driven Reconstruction of F(z)
The comoving sound horizon $\rm{r_{s}}$ at the drag epoch can be related to the comoving sound horizon $\rm{r_{*}}$ at the redshift of recombination ($\rm{z_{*}=1090}$) by $\rm{r_{d}\approx 1.02r_{*}}$. The quantity $r_{*}$ is inferred from CMB observation, using the position of the first peak in the CMB temperature power spectrum $\rm{\theta_{*}=r_{*}/(1+z_{*})D^{EM}_{A}(z_{*})}$ (Spergel et al., 2007; Hinshaw et al., 2013; Aghanim et al., 2020). As a result, the product of all these observable quantities will lead to an identity value at all redshifts if GR and the fiducial $\Lambda$CDM model of cosmology is the correct theory. However, if there is any departure from these models, then we can measure a deviation from one with redshift. It is important to note that the value of $\rm{r_{s}}$ is constant in the redshift range relevant for this study ( $\rm{z<2}$). So, even if there is an inaccuracy in the inference of true $\rm{r_{s}}$, the value of $\rm{\mathcal{F}(z)}$ will vary by an overall normalization. But the reconstructed redshift evolution of $\rm{\mathcal{F}(z)}$ will remain the same. In one of the later sections, we will show results for both the cases, (i) only $\rm{\mathcal{F}(z)}$ and (ii) $\rm{\mathcal{F}(z)}$ and $\rm{H_{0}}$ together. The second case will make it possible to marginalise over any inaccurate inference of sound horizon $\rm{r_{s}}$ due to incorrect inference of the Hubble constant. In the above equation, $\rm{D_{l}^{GW}(z)}$ is measured from GW sources, $\rm{\theta_{BAO}}$ is measured from galaxy two point correlation function, and $\rm{r_{s}}$ is inferred from the CMB temperature fluctuations. The redshift of the source can be inferred using cross-correlation of galaxy surveys with dark sirens. In Subsection 4, we provide a detailed framework for inferring redshift using cross-correlation with redshift-known galaxies.
@@ 3.1 Modelling of the astrophysical population of GW source
Merger rate: In this study we use a delay time model of the binary mergers (O’shaughnessy et al., 2010; Dominik et al., 2015a). In the delay time model, the merger rate is described in terms of the delay time distribution, denoted as $\rm{t_{d}}$. The delay time refers to the elapsed time between the formation of stars that will eventually become black holes and the actual merging of these black holes. It is important to note that the time delay is not uniform across all binary black holes but instead follows a specific distribution. This distribution function accounts for the variations in the delay time and is defined as follows
@@
where $T_{obs}$ indicates the total observation time, $\rm{\frac{dV_{c}}{dz}}$ corresponds to the comoving volume element, and $\rm{R(z)}$ denotes the merger rate (Karathanasis et al., 2022). We consider the delay time merger rate with a specific minimum delay time $\rm{t_{d}=500\text{Myrs}}$ and a power-law exponent of $\rm{d=1}$(Banerjee et al., 2010; Dominik et al., 2015b; Cao et al., 2018). To determine which events are detectable, the calculation of the matched filtering signal-to-noise ratio (SNR) plays a crucial role. The SNR serves as a measure of the strength of the GW signal relative to the background noise. Only those events with a matched filtering SNR greater than or equal to a predetermined threshold SNR ($\rm{\rho_{\rm TH}}$) can be reliably detected. (Maggiore, 2007).
@@
In this study, we employ two sets of GW detectors: one comprising LIGO (Aasi et al., 2015), Virgo (Acernese et al., 2014), and KAGRA (Akutsu et al., 2021) (LVK), and the other consisting of CE with two different configurations: a 20 km configuration (CE(20Km)+ET) and a 40 km configuration (CE(40Km)+ET), along with ET (Punturo et al., 2010). We establish different SNR thresholds for different GW observatories: $\rm{\rho_{\rm Th}=12}$ for the LVK (O5 sensitivity) system and $\rm{\rho_{\rm Th}=20}$ for both configurations of the CE+ET. We adopt a redshift bin size of $\rm{\Delta z=0.05}$, within which the total number of events is computed using Equation (9). Event generation is based on distances derived from redshifts, parameter $\rm{\Theta}$, and the mass distributions of the objects. Over a five-year operational period with a 75% duty cycle, the expected number of detectable events is projected to be approximately $\sim$13000 for LVK, $\sim$177000 for CE(40Km)+ET, and $\sim$144000 for CE(20Km)+ET, assuming a local merger rate ($\rm{R_{0}}$) of 20 $\rm{Gpc^{-3}yr^{-1}}$. It is important to note that for this analysis, we have specifically utilized the CE(20 km)+ET configuration. From now onwards, this combination is defined as CEET.
@@
In Figure 2, we present the projected total and detectable BBH events within specific redshift intervals ($\rm{\Delta z=0.05}$). To obtain these estimates, we first calculate the number of BBH mergers for each redshift bin using Equation (9). For each BBH event, we use the inverse transform method to determine the mass of the two black holes and the parameter $\rm{\Theta}$, drawing from their respective probability distributions. The masses of the black holes are generated within the range of $\rm{5M_{\odot}}$ to $\rm{50M_{\odot}}$, and $\rm{\Theta}$ is considered within the interval from 0 to 4. After incorporating the redshift information for each event, we calculate the corresponding luminosity distance and the SNR using Equation (14) for a single detector. The total SNR, denoted as $\rm{\rho_{total}}$, is calculated by combining the individual SNRs from the network of detectors using the formula $\rm{\rho_{total}=\sqrt{\sum_{i}\rho_{i}^{2}}}$. This analysis considers the detection capabilities of both the CEET and the LVK network over a specified observation period of 5 years with a 75% duty cycle. The curve depicted in the graph illustrates the distribution of these events, providing insights into the temporal occurrence of these mergers through cosmic history and their detection likelihood with current and future GW observatories.
@@ 4 Cross-Correlation Technique for Redshift Estimation of Dark Sirens
where $\rm{\delta^{r}_{\text{GW}}(k,z)}$ is the original density field of GW sources without sky localization impact, and $\rm{k_{\text{eff}}(z)}$ denotes a characteristic wavenumber varying with redshift $\rm{z}$ (Mukherjee et al., 2021a, b). $\rm{k_{\text{eff}}(z)}$ is the comoving scale defined as
@@
where $\rm{D_{c}(z)}$ is the comoving distance at redshift $\rm{z}$ (Mukherjee et al., 2021a). The exponential term $\rm{e^{-\frac{k^{2}}{k_{\text{eff}}^{2}(z)}}}$ quantifies the smoothing effect caused by the sky localization error, with a rapid decrease beyond $\rm{k>k_{\text{eff}}}$. Ongoing GW detectors like LVK are helping us characterize detected GW sources. In the future, upcoming ground-based detectors such as CE, ET, and LIGO-India, along with space-based detectors like LISA, will enhance our ability to characterize these sources. These advancements will improve our measurements of the redshifts of GW sources, thereby providing better insights into the Universe’s large scale structure and evolution.
In the standard model of cosmology, galaxies trace the underlying distribution of dark matter through the galaxy bias parameter, $\rm{b_{g}(k,z)}$. Consequently, GW sources are also expected to trace the dark matter distribution, but with their own distinct bias parameter, $\rm{b_{GW}(k,z)}$, which differs from the galaxy bias parameter.
The spatial clustering of galaxies and GW sources can also be described in Fourier space using the three-dimensional auto-power spectrum and cross-power spectrum at different redshifts $\rm{z}$. The matter distribution exhibits a clustering property that can be statistically described by the correlation function $\rm{\xi(r)}$, which is the fourier transform of the matter power spectrum $\rm{P_{m}(k)=\langle\delta_{m}(k)\delta_{m}(k)\rangle}$. Under the standard cosmological model, both GW events and the distribution of galaxies are biased tracers of the underlying matter distribution. Therefore, the cross-correlation of the power spectrum between GW sources and galaxy surveys can be used to infer the host redshift shell of the GW sources. For a detailed mathematical formulation, refer to Subsection 4.1.